OpenStudy (anonymous):

An isosceles triangle has base B and height H. Assuming its area remains constant, express the rate of change of the base with respect to the height.

OpenStudy (anonymous):

I'm getting -2A/h^2 Does that look right?

OpenStudy (anonymous):

My reasoning: A= 1/2 b * h 2A = b * h b = 2A/h db/dh 2A/h = db/dh 2A * h^-1 = 2A * -h^-2 = -2A/h^2

OpenStudy (anonymous):

Argh! Distributed wrongly.

OpenStudy (anonymous):

Right, but you don't differentiate Area...only height.

OpenStudy (anonymous):

Or am I wrong about this?

OpenStudy (anonymous):

Doesn't it mean that if we keep the Area constant, at what rate does the base change respective of the height? That can't be 0, I wouldn't think.

OpenStudy (anonymous):

A solution using Mathematica is attached.

OpenStudy (anonymous):

I think I made a mistake.

OpenStudy (anonymous):

yes, -2A/h^2 is the correct answer.

OpenStudy (anonymous):

Thanks very much!